# Impact of Forecast Stability on Navigational Contrail Avoidance

### v1

@@ 3.1 Forecast errors and the minimum segment length constraint
Proximity distributions suggest that these low ETS values are likely linked to the double-penalty problem. That is, they suggest that frequent pointwise errors are the result of relatively small displacement errors in positions of localized contrail-forming regions. Waypoints in ERA5 ISSRs and high-EF regions are statistically likely to be close (within 1 hour’s flight time) to ISSRs and high-EF regions in HRES forecasts, and waypoints in IAGOS ISSRs are similarly likely to be close to an ERA5 ISSR (Figures 4(a)-4(b), solid lines). This proximity is not merely the result of most waypoints being close to contrail-forming regions: waypoints outside HRES and ERA5 contrail-forming regions are typically much farther from contrail-forming regions in ERA5 and IAGOS measurements, respectively (Figures 4(a)-4(b), dashed lines). The proximity between HRES and ERA5 contrail-forming regions has some dependence on lead time: the fraction of waypoints in ERA5 ISSRs that are within 1 hour’s flight time of an HRES ISSR decreases from above 95% at short lead times to slightly above 90% at 48 hour lead times, and the equivalent statistic for high-EF regions varies from slightly above 90% to around 85% (Figure 4(c)).
Figure 4: (a) Proximity distributions for HRES vs. ERA5 ISSRs (blue) and ERA5 vs IAGOS ISSRs (black). (b) Proximity distributions for HRES vs ERA5 high-EF regions. (c) Fraction of waypoints in ERA5 ISSRs that are within an hour’s flight time of an HRES ISSR (blue) and of waypoints in ERA5 high-EF regions that are within an hour’s flight time of an HRES high-EF region (red), binned by forecast lead time. High EF regions are defined as regions with EF per unit flight distance greater than $10^{7}$ J m-1. Flights with a duration less than four hours are excluded to avoid large differences in the average duration of optimized trajectories and IAGOS flights. In panel (c), “HRES positives” refer to ISSRs and high-EF regions for blue and red points, respectively.
@@ 4 Forecast stability experiments
We used the contrail forecasts described in Section 2 and the optimizer described in Section 3 to evaluate the impact of lead-time-dependent forecast errors on the effectiveness of navigational avoidance. We first selected one random day per month during each month of 2019, and then selected 5000 random commercial flights with takeoff times within the 24 hour period starting at 01Z on each selected day. We chose to select one random day per month—rather than 12 random days throughout 2019—to ensure that we sampled a wide range of meteorological regimes throughout a full seasonal cycle. We obtained aircraft trajectory information for selected flights from automatic dependent surveillance-broadcast (ADS-B) telemetry purchased from Spire Aviation.
@@
Optimizing trajectories using a shorter minimum segment length decreases the fraction of cost-optimal contrail EF avoided by contrail-optimal trajectories (Figure 6(b)), indicating that a sufficiently long minimum segment length helps to limit the impact of forecast errors. The effect of varying the minimum segment length is particularly pronounced at the longest lead times considered in this study. At 42-48 hour lead times, decreasing the minimum segment length from 90 to 10 minutes leads to a drop in relative EF reduction from above 70% to below 60%. The effect is somewhat more modest at the 8-24 hour lead times most relevant for flight planning, but even here reducing the minimum segment length leads to $\sim$10decreases in the fraction of contrail EF avoided. We note that a longer minimum segment length also has operational utility in that trajectories with less frequent climbs and descents reduce workload for pilots and air traffic control, and also reduce maintenance costs due to engine wear.
@@ 5 Discussion
HRES forecasts with 8-24 hour lead times and ERA5 reanalysis frequently disagree on the precise locations where high-warming contrails can form (recall Figure 3). Nevertheless, our results indicate that trajectory optimization based on HRES forecasts can lead to large (70-100%) reductions in contrail forcing evaluated using ERA5 reanalysis, particularly when using an optimizer with a minimum segment length constraint tuned to take advantage of spatial proximity of HRES and ERA5 contrail-forming regions (recall Figure 4 and 5(a)). Here, we examine in more detail how our optimzer achieves large EF reductions. While the following discussion focuses on results from our optimizer, it provides a framework that could be used to analyze how the performance of other optimizers is affected by spatiotemporal errors in contrail forecasts.
@@
In this discussion, we show that our optimizer’s minimum segment length constraint aids in accomplishing (i) by increasing the probability that waypoints close to but not within forecast high-EF regions are included in avoidance maneuvers (Section 5.1), and we show that our optimizer accomplishes (ii) largely by deviating downward to levels that are too warm for contrail formation (Section 5.2). We conclude by discussing the potential for a similar optimizer to effectively accommodate errors in forecasts relative to reality rather than relative to reanalysis.
@@ 5.3 Implications for assessing forecast accuracy
In Section 3.1, we showed that pointwise error rates in predictions of ISSR locations are similar in forecasts compared to reanalysis and reanalysis compared to reality (Figure 3). We also showed that proximity distributions, which characterize the spatial structure of forecast errors, are remarkably similar when comparing forecasts to reanalysis and reanalysis to IAGOS measurements. In particular, nearly all ($>90\%$) of waypoints inside IAGOS ISSRs are within an hour’s flight time of an ERA5 ISSR, just as nearly all waypoints inside ERA5 ISSRs are within an hour’s flight time of ISSRs in HRES forecasts. This indicates that errors in ISSR locations in forecasts relative to reanalysis and reanalysis relative to reality occur at similarly-small spatial scales, and suggests that an optimizer that accommodates small-scale errors in forecasts relative to reanalysis may also be capable of accommodating errors in forecasts relative to reality.
Our analysis of the effectiveness of avoidance maneuvers (Section 5.2) requires information about pairs of cost- and contrail-optimal trajectories and so cannot easily be adapted to use measurements from IAGOS flights. However, one key result—that avoidance maneuvers targeting formation are highly effective when considering errors between forecasts and reanalysis—may remain true when considering errors between forecasts and the real world. This is because forecasts of UTLS temperature are better-constrained by observations than forecasts of UTLS humidity and show relatively good agreement with IAGOS measurements [29, 44], and contrail formation is primarily temperature- rather than humidity-dependent. Further, recent experiments and modeling suggests that higher ambient temperatures significantly reduce contrail impacts from lean-burn engines [21]. This suggests that updated contrail models that more accurately account for vPM activation may in turn exhibit increased forecast skill since temperature is both a stable and accurate forecast feature.
@@ 5.4 Operational Implications
We note that in practice, navigational avoidance of persistent contrails will involve additional fuel burn and changes in flight times. In A, we present a detailed discussion of the penalties associated with the avoidance profiles used in this study. Specifically, avoidance of approximately 90% of contrail energy forcing is achieved with an approximate fuel penalty of 0.3% additional fuel burn. While this number is not insignificant, we note that more intelligent optimization procedures (e.g., strategies using lateral as well as vertical avoidance) may substantially reduce this penalty (see A for futher discussion). Other operational considerations, such as effects on workload, airspace congestion, interactions with other meteorological safety considerations (e.g., turbulence and icing), are important but left to future work.
@@ 6 Conclusion
Our results suggest that forecasts of contrail EF at lead times relevant to flight planning ($<48\,\mathrm{hrs}$) are sufficiently stable to support navigational contrail avoidance in the flight planning stage. Contrail-optimal trajectories generated based on HRES forecasts with lead times of up to 6 hours reduce contrail radiative forcing assessed using ERA5 reanalysis by amounts exceeding 90%. Trajectory optimization using forecasts with longer lead times most relevant for flight planning (8-24 hrs) is slightly less effective, but can still reduce reanalysis-assessed EF by 80-90%. Our methods—trajectory optimization that attempts to reduce contrail warming using forecasts generated with a simplified contrail process model, followed by a “retrospective” assessment using reanalysis meteorology and a high-fidelity contrail process model—intentionally mimic a workflow that could be used by operators to manage and report contrail impacts. Our results demonstrate that, within such a workflow, the reported reduction in contrail warming is unlikely to be strongly sensitive to the lead time of the forecast used for flight planning.
@@
Finally, our work assumes that flights follow optimized trajectories exactly. This will not be true in practice: delays or other operational issues (e.g., the need to avoid turbulence, or traffic-related restrictions) may cause aircraft to deviate from contrail-optimal flight plans. In practice, implementation of navigational avoidance will necessitate some amount of tactical decision making by pilots, dispatchers, and air-traffic managers. The impact of these deviations on contrail EF has not been quantified here but will be a focus of future studies and operational trials.
@@ Appendix A Fuel and Delay Penalties
Figure 9: (a) The empirical cumulative distribution of changes in flight times for contrail avoidance flights. The majority ($>58$%) of avoidance flights are of equal or shorter in duration due to the fact that they almost always fly at lower altitudes with higher true air speeds. (b) Fleet-wide fuel penalties for different contrail cost indexes and the associated contrail energy forcing achieved with each. Here, $\gamma$ represents a scaling factor applied to the value of 6.7e-11 kg/J used for the results in the main body of the paper.
@@
Example fuel burn penalties are depicted in Figure 9(b). This figure depicts the trade-off between excess fuel burn and contrail EF reduction, obtained by adjusting the contrail cost index. Here, $\gamma$ is used as a scaling factor applied the value of 6.7e-11 kg/J used in the results presented in the main body of this paper; that is, $\gamma=1$ indicates a contrail cost index of 6.7e-11, and $\gamma=0$ indicates the cost-optimal route. The values in Figure 9(b) were obtained by optimizing flights against only the forecast from 12Z the day before flight departure (lead-times of 12-36 hours) rather than the entire set of forecasts used previously. We note that fuel penalties associated with avoidance strategies are unlikely to be sensitive to forecast lead times.
@@
By examining Figure 9(b), we see that vertical deviations can produce a 86% EF reduction with 0.33% additional fuel burn. The near vertical slope of the line at the right-hand side of the figure suggests that further increases to the contrail cost index would not achieve significant additional EF reduction. On the left-side side of the figure, it can be seen that over 33% EF reduction can be achieved with less than 0.01% additional fuel burn if the cost index is decreased by a factor of 1000. As stated previously, we believe a more sophisticated optimizer (particularly one that considers lateral as well as vertical deviations) may be able to achieve similar EF reductions with a smaller fuel penalty. This is a topic of forthcoming research.

### v3

@@ 3.1 Forecast errors and the minimum segment length constraint
Low ETS scores suggest a large number of pointwise errors between the two products. However, if these pointwise errors are the result of relatively small displacement errors in positions of localized contrail-forming regions, then–just as failing to predict the precise location of a thunderstorm is not necessarily a barrier to deciding to issue a county-level tornado watch–failing to predict the precise locations of contrail-forming regions may not be a barrier to identifying volumes of airspace that flights should avoid. Indeed, “proximity distributions” of distances between between forecast, reanalysis, and observed contrail-forming regions (Figure 4(a)-4(b)) suggest that displacement errors are usually small. Waypoints in ERA5 ISSRs and high-EF regions are statistically likely to be close (within 1 hour’s flight time) to ISSRs and high-EF regions in HRES forecasts, and waypoints in IAGOS ISSRs are similarly likely to be close to an ERA5 ISSR (Figures 4(a)-4(b), solid lines). This proximity is not merely the result of most waypoints being close to contrail-forming regions: waypoints outside HRES and ERA5 contrail-forming regions are typically much farther from contrail-forming regions in ERA5 and IAGOS measurements, respectively (Figures 4(a)-4(b), dashed lines). The proximity between HRES and ERA5 contrail-forming regions has some dependence on lead time: the fraction of waypoints in ERA5 ISSRs that are within 1 hour’s flight time of an HRES ISSR decreases from above 95% at short lead times to slightly above 90% at 48 hour lead times, and the equivalent statistic for high-EF regions varies from slightly above 90% to around 85% (Figure 4(c)).
Figure 4: (a) Cumulative distribution functions (CDFs) measuring the distance between HRES vs. ERA5 ISSRs (blue) and ERA5 vs IAGOS ISSRs (black), henceforth referred to as proximity distributions. (b) Proximity distributions for HRES vs ERA5 high-EF regions. (c) Fraction of waypoints in ERA5 ISSRs that are within an hour’s flight time of an HRES ISSR (blue) and of waypoints in ERA5 high-EF regions that are within an hour’s flight time of an HRES high-EF region (red), binned by forecast lead time. High EF regions are defined as regions with EF per unit flight distance greater than $10^{7}$ J m-1. Flights with a duration less than four hours are excluded to avoid large differences in the average duration of optimized trajectories and IAGOS flights. In panel (c), “HRES positives” refer to ISSRs and high-EF regions for blue and red points, respectively.
@@ 4 Forecast stability experiments
We used the contrail forecasts described in Section 2 and the optimizer described in Section 3 to evaluate the impact of lead time-dependent forecast errors on the effectiveness of navigational avoidance. We first selected one random day per month during each month of 2019, and then selected 5000 random commercial flights with takeoff times within the 24 hour period starting at 01Z on each selected day. We chose to select one random day per month—rather than 12 random days throughout 2019—to ensure that we sampled a wide range of meteorological regimes throughout a full seasonal cycle. In B, we discuss month trends and present evidence to suggest that a larger sample of days is unlikely to change the results presented in this section. We obtained aircraft trajectory information for selected flights from automatic dependent surveillance-broadcast (ADS-B) telemetry purchased from Spire Aviation.
@@
Optimizing trajectories using a shorter minimum segment length decreases the fraction of cost-optimal contrail EF avoided by contrail-optimal trajectories (Figure 6(b)), indicating that a sufficiently long minimum segment length helps to limit the impact of forecast errors. The effect of varying the minimum segment length is particularly pronounced at the longest lead times considered in this study. At 42-48 hour lead times, decreasing the minimum segment length from 90 to 10 minutes leads to a drop in relative EF reduction from above 70% to below 60%. The effect is somewhat more modest at the 8-24 hour lead times most relevant for flight planning, but even here reducing the minimum segment length leads to $\sim$10% decreases in the fraction of contrail EF avoided. We note that a longer minimum segment length also has operational utility in that trajectories with less frequent climbs and descents reduce workload for pilots and air traffic control, and also reduce maintenance costs due to engine wear.
@@ 5 Discussion
HRES forecasts with 8-24 hour lead times and ERA5 reanalysis exhibit a poor pointwise agreement on the precise locations where high-warming contrails can form (ETS << 0.4, recall Figure 3). Despite this, we show that trajectory optimization based on HRES forecasts can lead to large (70-100%) reductions in contrail forcing evaluated using ERA5 reanalysis, particularly when using an optimizer with a minimum segment length constraint tuned to take advantage of spatial proximity of HRES and ERA5 contrail-forming regions (recall Figure 4 and 5(a)). Here, we examine in more detail how our optimzer achieves large EF reductions. While the following discussion focuses on results from our optimizer, it provides a framework that could be used to analyze how the performance of other optimizers is affected by spatiotemporal errors in contrail forecasts.
@@
In this discussion, we show that our optimizer’s minimum segment length constraint aids in accomplishing (i) by increasing the probability that waypoints close to but not within forecast high-EF regions are included in avoidance maneuvers (Section 5.1), and we show that our optimizer accomplishes (ii) largely by deviating downward to levels that are too warm for contrail formation (Section 5.2). We conclude by discussing the potential for a similar optimizer to effectively accommodate errors in forecasts relative to reality rather than relative to reanalysis (Section 5.3).
@@ 5.3 Implications for assessing forecast accuracy
In Section 3.1, we showed that pointwise error rates in predictions of ISSR locations are similar in forecasts compared to reanalysis and reanalysis compared to reality (Figure 3). We also showed that proximity distributions, which characterize the spatial structure of forecast errors, are remarkably similar when comparing forecasts to reanalysis and reanalysis to IAGOS measurements. In particular, nearly all ($90\%$) of waypoints inside IAGOS ISSRs are within an hour’s flight time (either forward or backward along the current trajectory) of an ERA5 ISSR at the same altitude found in the aircraft trajectory, just as nearly all waypoints inside ERA5 ISSRs are within an hour’s flight time of ISSRs in HRES forecasts. This indicates that errors in ISSR locations in forecasts relative to reanalysis and reanalysis relative to reality occur at similarly-small spatial scales, and suggests that an optimizer that accommodates small-scale errors in forecasts relative to reanalysis may also be capable of accommodating errors in forecasts relative to reality.
Our analysis of the effectiveness of avoidance maneuvers (Section 5.2) requires information about pairs of cost- and contrail-optimal trajectories and so cannot easily be adapted to use measurements from IAGOS flights. However, one key result—that avoidance maneuvers that avoid formation are likely to be successful when considering errors between forecasts and reanalysis—may remain true when considering errors between forecasts and the real world. This is because forecasts of UTLS temperature are better-constrained by observations than forecasts of UTLS humidity and show relatively good agreement with IAGOS measurements [32, 47], and contrail formation is primarily temperature- rather than humidity-dependent. Further, recent experiments and modeling suggests that higher ambient temperatures significantly reduce contrail impacts from lean-burn engines [21]. This suggests that updated contrail models that more accurately account for vPM activation may in turn exhibit increased forecast skill since temperature is both a stable and accurate forecast feature. It is worth noting that these findings do not imply that performing navigational avoidance based on formation criteria alone would be advisable, as such a strategy would result in prohibitively large fuel penalties.
@@ 5.4 Operational Implications
We note that in practice, navigational avoidance of persistent contrails will involve additional fuel burn and changes in flight times. In A, we present a detailed discussion of the penalties associated with the avoidance profiles used in this study. Specifically, avoidance of approximately 80%-90% of contrail energy forcing is achieved with an approximate fuel penalty between 0.2% and 0.3% additional fuel burn depending on the choice of minimum segment length. While this number is not insignificant, we note that more intelligent optimization procedures (e.g., strategies using lateral as well as vertical avoidance) may substantially reduce this penalty (see A for futher discussion). Other operational considerations, such as effects on workload, airspace congestion, interactions with other meteorological safety considerations (e.g., turbulence and icing), are important but left to future work.
Finally, we note that we have performed our simulations with a single cost index setting. In practice, operators plan trajectories at a variety of cost indices based on various economic and operational considerations. An important future piece of work would be to perform a range of contrail avoidance simulations at varying cost indices to understand the effects of variable cost indices on the cost penalties and efficacy of navigational contrail avoidance. Such a study may be best carried out by a commercial grade flight planning system similar to [16].
@@ 6 Conclusion
Our results suggest that forecasts of contrail EF at lead times relevant to flight planning ($48\,\mathrm{hrs}$) are sufficiently stable to support navigational contrail avoidance in the flight planning stage. Contrail-optimal trajectories generated based on HRES forecasts with lead times of up to 6 hours reduce contrail radiative forcing assessed using ERA5 reanalysis by amounts exceeding 90%. Trajectory optimization using forecasts with longer lead times most relevant for flight planning (8-24 hrs) is slightly less effective, but can still reduce reanalysis-assessed EF by 80-90%. Our methods—trajectory optimization that attempts to reduce contrail warming using forecasts generated with a simplified contrail process model, followed by a “retrospective” assessment using reanalysis meteorology and a high-fidelity contrail process model—intentionally mimic a workflow that could be used by operators to manage and report contrail impacts. Our results demonstrate that, within such a workflow, the reported reduction in contrail warming is unlikely to be strongly sensitive to the lead time of the forecast used for flight planning.
@@
Finally, our work assumes that flights follow optimized trajectories exactly. This will not be true in practice: delays or other operational issues (e.g., the need to avoid turbulence, or traffic-related restrictions) may cause aircraft to deviate from contrail-optimal flight plans. Additional factors, such workload considerations by pilots or dispatchers, or other human factors, may further limit the efficacy of navigational avoidance. In practice, implementation of navigational avoidance will necessitate some amount of tactical decision making by pilots, dispatchers, and air-traffic managers. The impact of these deviations on contrail EF has not been quantified here but will be a focus of future studies and operational trials.
@@ Appendix A Fuel and Delay Penalties
Figure 9: (a) The empirical cumulative distribution of changes in flight times for contrail avoidance flights. The majority (>58>58%) of avoidance flights are of equal or shorter in duration due to the fact that they almost always fly at lower altitudes with higher true air speeds. (b) Fleet-wide fuel penalties for different contrail cost indexes and the associated contrail energy forcing achieved with each. Here, $\gamma$ represents a scaling factor applied to the value of 6.7e-11 kg/J used for the results in the main body of the paper. (c) Fleet-wise fuel penalties for different minimum segment lengths considering cost optimal ($\gamma=0$), and contrail optimal ($\gamma=1$). Imposing a 90-minute minimum segment length constraint produces a fuel burn penalty between 0.05% and 0.1% compared to a baseline obtained using the $\gamma=0$ and a 10-minute minimum segment length.
@@
Example fuel burn penalties are depicted in Figure 9(b). This figure depicts the trade-off between excess fuel burn and contrail EF reduction, obtained by adjusting the contrail cost index. Here, $\gamma$ is used as a scaling factor applied the value of 6.7e-11 kg/J used in the results presented in the main body of this paper; that is, $\gamma=1$ indicates a contrail cost index of 6.7e-11, and $\gamma=0$ indicates the cost-optimal route. The values in Figure 9(b) were obtained by optimizing flights against only the forecast from 12Z the day before flight departure (lead times of 12-36 hours) rather than the entire set of forecasts used previously. We note that fuel penalties associated with avoidance strategies are unlikely to be sensitive to forecast lead times.
@@
By examining Figure 9(b), we see that vertical deviations can produce a 86% EF reduction with 0.33% additional fuel burn. The near vertical slope of the line at the right-hand side of the figure suggests that further increases to the contrail cost index would not achieve significant additional EF reduction. On the left-side side of the figure, it can be seen that over 33% EF reduction can be achieved with less than 0.01% additional fuel burn if the cost index is decreased by a factor of 1000.
In Figure 9(c) we examine the fuel burn penalty caused by the minimum segment length constraint of the optimizer. We note that for nearly all trajectories in this study, the spacing used in the trajectory optimization search procedure has a segment length greater than 10 minutes, and so the 10-minute minimum segment length constraint may be interpreted as no constraint. In this figure, we see that at a contrail cost index of 0, the 90-minute constraint produces approximately a 0.05% fuel burn penalty, and at a contrail cost index of 1, we see approximately a 0.1% fuel burn penalty compared to the case of the 10-minute minimum segment length. We note that the 10-minute minimum segment length constraint may be unrealistically short for use in commercial flight planning, and that commercial-grade optimizers typically constrain the number of step-climbs by imposing engine-cycle costs which penalize changes in thrust, rather than constraining the time between step climbs.
We again re-iterate that we provide the fuel burn estimates here as very rough approximations to the cost of contrail avoidance, and highlight that the optimization procedure presented in this work is highly flexible. As stated previously, we believe a more sophisticated optimizer (particularly one that considers lateral as well as vertical deviations) may be able to achieve similar EF reductions with a smaller fuel penalty. This is a topic of forthcoming research.
Appendix B Monthly Variability
Figure 10: (a) The total fleet-wide contrail energy forcing for each day sampled in the study. One day per month was chosen at random and 5000 flights were simulated on each day. The impact of the cost optimal trajectories is shown in orange which exhibits expected seasonal variability. The resulting contrail impact after optimization according to a short-range forecast is shown in blue. (b) The percentage of contrail warming removed by month, with 95% confidence intervals. Apart from one outlier, the results are well clustered around the mean, and all months obtain greater than 80% reduction.
Figure 11: Change in fleet-wide contrail EF as a function of forecast lead time by month. While there is some spread in total reduction by month, all months follow the same overall trend of forecast stability, with a reduction of mitigated EF of 10%-20% as lead times increase from 1 to 48 hours
In this appendix, we consider monthly trends of contrail warming reduction. In this study, we sampled one day per month, and for each day simulated a total of 5000 flights. We chose to sample one day per month at random rather that a random sampling of 12 days in order to sample a wide range of meteorological regimes. In Figure 10(a) we show the total fleet-wide contrail impact for each day sampled in the study. Here, we depict both the impact of the cost optimal trajectories, along with the residual contrail impacts for flights optimized to mitigate contrail impacts. The flights depicted in this figure were optimized with a minimum segment length of 90 minutes using a short-range forecast with a maximum lead time of 24 hours. Contrail impacts were evaluated using ERA5 reanalysis and the trajectory version of CoCiP. Figure 10(b) shows the same data, but depicts the relative EF reduction between the two sets of flight trajectories. The error bars depict the 95% confidence interval obtained by bootstrapping per-flight. The results are well clustered around the sample mean of 90.5% [89.4%, 91.7%], with no month achieving an EF reduction below 80%. We note that in outlier month of August, the large fractional reduction in contrail EF does not much change the 12 month average results because of the relatively small amount of EF avoided in August.
In Figure 11, we show the EF reduction achieved when binning the results by forecast lead time. This figure shows the same data depicted in Figure 6(a) for each individual month in the study. While there is some variability in the total mitigation rate by month, as previously depicted in Figure 10(b), the overall trend by lead time for each month shows very little variability. For all months, the total reduction in EF at 36 hour lead times is less than 20% lower than the reduction at short lead times. We note that there is slightly more variability in the amount of mitigation obtained with lead times longer than 36 hours, with some, but not all, months exhibiting further degradation.
These analyses suggest that the central result claimed in this work—that, compared to reanalysis, 80%-90% of contrail warming may be mitigated through the flight planning using forecasts with lead times of 8-24 hours—is unlikely to change by sampling additional days.